OpenStudy (anonymous):

The following statement is a correct argumentation: 13 is prime, since divisible by 3 21 is prime, since divisible by 3 23 is prime, since not divisible by 3 25 is not prime, since divisible by 5 7 is prime, since not divisible by 2, 3, 4, 5 and 6

OpenStudy (anonymous):

False True True True True

OpenStudy (anonymous):

Wow sorry the second one is False.

OpenStudy (anonymous):

3 is false. 23 is prime, but not because its not divisible by 3. It's because it's not divisible by any integer before it except 1.

OpenStudy (anonymous):

I agree Xavier, but the statement depending on how you take it is true.

OpenStudy (anonymous):

I'm taking the format as: Statement, Reason.

OpenStudy (anonymous):

I can agree then. 3 would be false.

OpenStudy (anonymous):

The following product is a complete decomposition into prime factors: 12 = 2 × 2 × 3 14 = 2 × 7 16 = 4 × 4 42 = 6 × 7 42 = 2 × 3 × 7 63 = 3 × 21

OpenStudy (anonymous):

Yes gthug that is correct

OpenStudy (anonymous):

Well 16 woudl be 2 * 2 * 2 * 2

OpenStudy (anonymous):

63 would be 3 * 3 * 7